OpenStudy (anonymous):

Your supplier only ships acid solution in 7% and 15% concentrations. How many liters of the 7% solution should you mix to obtain 8 liters of 10% solution?

OpenStudy (anonymous):

You need to mix 5 liters of the 7% solution. :'D

OpenStudy (amistre64):

acid@7% = 8(.10-.15)/(.07-.15)

OpenStudy (amistre64):

5 is good :)

OpenStudy (amistre64):

assuming of course that your mixing the 2 together

OpenStudy (anonymous):

If you need the equation I used, it's 7x + 15(8-x) = 80 : )

OpenStudy (amistre64):

oh i used it to; but more condensified ;)

OpenStudy (anonymous):

Lol, yes. :D

OpenStudy (anonymous):

how can i write this out. like show work

OpenStudy (amistre64):

depends on how you want to solve it ... usually the long way is to make equations that relate and either sub or eliminate to get to the answer

OpenStudy (amistre64):

a + b =t ; *-x ax + by = tz -ax -bx =-tx ax +by = tz ------------ b(y-x) = t(z-x) b @ y% = t(z-x)/(y-x)

OpenStudy (anonymous):

woah just the simple way please. using actual numbers

OpenStudy (anonymous):

Let x= liters of 7% sol'n 8-x = liters of 15% sol'n 7x + 15(8-x) = 10(8) Solve for x. : ) You can always use put the percents in decimal or fraction form, but I always find it easier to do it using whole numbers.

OpenStudy (amistre64):

lol ... a + b = 8 ;*-.15 a(.15) + b(.07) = 8(.10) -a(.15) + b(-.15) = 8(-.15) a(.15) + b(.07) = 8(.10) ------------------------ b(.07-.15) = 8(.10-.15) amount @ 7% = 8(.10-.15)/(.07-.15)